# Systematic Classification of History-Dependent Invariants in the Lorenz System: Regularization Classes and Dynamical Signatures

### v1

@@ III.3 Structural Origin of the Regularized Evolution Equation
A key technical point deserves explicit attention: why does the evolution equation for the regularized variable $v_{n}$ emerge as a function of $(x,y,z)$ alone, independent of $v_{n}$ itself? This property is not assumed but follows structurally from the construction.
Consider an invariant of the general form $K_{n}=P_{n}(x,y,z)+p(x,y,z)\cdot u_{n}$. The conservation requirement $dK_{n}/dt=0$ yields:
@@
Solving for $\dot{u}_{n}$ (assuming $p(x,y,z)\neq 0$; the behavior at nullclines where $p$ vanishes is addressed in Sec. III.4):
$\dot{u}_{n}=-\frac{1}{p}\left(\frac{dP_{n}}{dt}+\frac{dp}{dt}u_{n}\right)=-\frac{\dot{P}_{n}}{p}-\frac{\dot{p}}{p}u_{n}.$ (17)
The original variable $u_{n}$ has $u_{n}$-dependent dynamics. The term $(\dot{p}/p)u_{n}$ in Eq. (17) shows that $\dot{u}_{n}$ depends on $u_{n}$ itself. This means $u_{n}(t)$ satisfies a linear ODE with time-dependent coefficients, not a simple integral $u=\int f\,dt$. The $u_{n}$-dependent term is numerically stiff near nullclines where $p\to 0$.
The regularized variable $v_{n}$ has $v_{n}$-independent dynamics. Defining $v_{n}=p\cdot u_{n}$ and differentiating:
$\dot{v}_{n}=\dot{p}\cdot u_{n}+p\cdot\dot{u}_{n}=\dot{p}\cdot u_{n}+p\left(-\frac{\dot{P}_{n}}{p}-\frac{\dot{p}}{p}u_{n}\right)=-\dot{P}_{n}.$ (18)
The $u_{n}$-dependent terms cancel exactly, yielding $\dot{v}_{n}=Q_{n}(x,y,z)\equiv-\dot{P}_{n}$, which depends only on the physical coordinates. This algebraic cancellation is guaranteed by the structure of the invariant $K_{n}=P_{n}+v_{n}$: differentiating both sides of $K_{n}=c_{n}$ immediately gives $\dot{v}_{n}=-\dot{P}_{n}$ without reference to the intermediate variable $u_{n}$.
This distinction is essential: while the original auxiliary variable $u_{n}$ satisfies a stiff linear ODE, the regularized variable $v_{n}$ satisfies a simple integral $v_{n}(t)=v_{n}(0)+\int_{0}^{t}Q_{n}\,ds$. The regularized formulation is therefore preferred for both theoretical analysis and numerical implementation.
Sign convention for Class III. The derivation above assumes $v=p\cdot u$. For Class III, we adopt the convention $v_{n}^{\mathrm{III}}=-p_{\mathrm{III}}\cdot u_{n}$ (Eq. 14). This sign flip does not affect the fundamental result $\dot{v}_{n}=-\dot{P}_{n}$, which follows directly from differentiating $K_{n}=P_{n}+v_{n}=c_{n}$ regardless of how $v_{n}$ is defined in terms of $u_{n}$. However, the evolution equation for the original auxiliary variable $u_{n}$ differs between classes. Since all analysis in this work employs the regularized variables $v_{n}$ exclusively, this sign convention affects only the formal relationship between $u_{n}$ and $v_{n}$, not the invariant calculations themselves.
Explicit verification for $K_{13}$. To confirm sign consistency, we verify directly that $K_{13}=P_{13}+v_{13}$ is conserved. With $P_{13}=y^{2}/2+xyz-xy\rho+x^{2}\sigma/2-xy\sigma$ and $v_{13}^{\mathrm{III}}=(\beta z-xy)u_{13}$, the invariant is:
$K_{13}=\frac{y^{2}}{2}+xyz-xy\rho+\frac{x^{2}\sigma}{2}-xy\sigma+v_{13}.$ (19)
Differentiating: $\dot{K}_{13}=\dot{P}_{13}+\dot{v}_{13}$. Setting $\dot{v}_{13}=-\dot{P}_{13}$ yields $\dot{K}_{13}=0$ as required. The evolution function $Q_{13}=\dot{v}_{13}=-\dot{P}_{13}$ exhibits large-amplitude spikes at separatrix crossings (as derived in Sec. VII.4), consistent with the intermittent behavior observed in Figure 2(c) and the elevated skewness magnitude $|S_{\mathrm{III}}|\approx 2.7$.
The cancellation fails for the null class (permutations 4abc) precisely because placing $u$ in the leading position violates the integrability conditions (Schwarz’s theorem) before the regularization step can be applied, as demonstrated in Theorem 1.
III.4 Singularity Structure and the Distinction Between $u_{n}$ and $v_{n}$
The regularization polynomials $p_{\mathrm{I}}$, $p_{\mathrm{II}}$, and $p_{\mathrm{III}}$ vanish on nullclines of the Lorenz flow. This has different consequences for $u_{n}$ and $v_{n}$.
The regularized variable $v_{n}$ is everywhere smooth and bounded. From conservation $K_{n}=P_{n}+v_{n}=c_{n}$:
$v_{n}(t)=c_{n}-P_{n}(x(t),y(t),z(t)).$ (20)
Since $P_{n}$ is polynomial and the attractor is compact, $v_{n}$ inherits bounded, smooth behavior. The evolution equation $\dot{v}_{n}=Q_{n}(x,y,z)$ involves only polynomial functions, ensuring $v_{n}$ varies continuously across nullclines.
The original auxiliary variable $u_{n}$ diverges at nullcline crossings. The relationship $v_{n}=p_{n}\cdot u_{n}$ implies
$u_{n}(t)=\frac{v_{n}(t)}{p_{n}(x(t),y(t),z(t))}=\frac{c_{n}-P_{n}(x,y,z)}{p_{n}(x,y,z)}.$ (21)
When the trajectory crosses a nullcline where $p_{n}=0$, $u_{n}\to\pm\infty$.
This divergence is a coordinate singularity, analogous to the singularity at $r=0$ in polar coordinates, not a physical singularity. The invariant $K_{n}=P_{n}+p_{n}u_{n}=P_{n}+v_{n}$ remains well-defined:
$\lim_{p_{n}\to 0}K_{n}=\lim_{p_{n}\to 0}\left[P_{n}+p_{n}\cdot\frac{c_{n}-P_{n}}{p_{n}}\right]=c_{n}.$ (22)
Proper formulation. The rigorous formulation employs regularized variables:
$K_{n}=P_{n}(x,y,z)+v_{n},\quad\dot{v}_{n}=Q_{n}(x,y,z).$ (23)
This representation is singularity-free throughout phase space.
III.5 Class I Invariants: Permutations 1abc
@@
III.6 Class II Invariants: Permutations 2abc
@@
III.7 Class III Invariants: Permutations 3abc
@@
III.8 Null Class: Permutations 4abc
The six permutations beginning with 4 yield no valid invariants.
Theorem 1 (Null Permutations).
Permutations 4abc yield no valid history-dependent invariants.
Proof.
We demonstrate for permutation 4123, corresponding to $(u,x,y,z)$. The orthogonality ansatz assigns
@@ III.8 Null Class: Permutations 4abc
Consider Schwarz’s theorem for mixed partial derivatives. From Eqs. (44) and (46):
@@
Equality $\partial_{yu}C=\partial_{uy}C$ requires $0=x$. This constraint restricts $C$ to the hyperplane $x=0$, which is not invariant under the Lorenz flow. No smooth invariant satisfying ansatz (42) exists throughout phase space.
For valid classes (1abc, 2abc, 3abc), the ansatz assigns $\partial_{u}C$ to $\dot{u}$, which is determined self-consistently. The flexibility to choose $\dot{u}$ allows the integrability conditions to be satisfied.
For null permutations 4abc, $\partial_{u}C$ is assigned to one of $\dot{x}$, $\dot{y}$, or $\dot{z}$, whose functional forms are prescribed by Eqs. (1)–(3). This eliminates the flexibility needed for consistency: the system becomes overdetermined.
Analogous contradictions arise for all permutations 4abc. Permutations 4213 and 4231 yield the constraint $0=y$; permutations 4312 and 4321 yield $0=\sigma(y-x)$; permutation 4132 yields $0=x$. Each constraint defines a lower-dimensional manifold not invariant under the flow. ∎
This result reflects violation of Schwarz integrability conditions when the auxiliary variable occupies the leading position. The ansatz generates an overdetermined system with no smooth solution throughout phase space.
@@ IV.1 Boundedness of Regularized Variables
A consequence of the conservation laws is that regularized variables $v_{n}(t)$ remain bounded for all time, despite their coupling to the trajectory history through $\dot{v}_{n}=Q_{n}(x,y,z)$.
From conservation $K_{n}=P_{n}(x,y,z)+v_{n}=c_{n}$:
@@
Since $P_{n}$ is polynomial and the Lorenz attractor is compact, $v_{n}(t)$ is bounded. The fluctuation range is determined by the variation of $P_{n}$ over the attractor.
Distinction from $u_{n}$. The original auxiliary variable $u_{n}(t)=v_{n}(t)/p_{n}(x,y,z)$ diverges at nullcline crossings (Sec. III.4). This coordinate singularity does not affect conservation. We work exclusively with regularized variables $v_{n}$.
@@
A nonzero average would imply unbounded growth of $v_{n}$, contradicting Eq. (49).
Numerical integration at standard parameters over $T=5000$ confirms these predictions with high precision.
IV.2 Structure of Pairwise Differences
Boundedness extends to differences of regularized variables. For invariants $K_{n}$ and $K_{m}$:
$v_{n}-v_{m}=(c_{n}-P_{n})-(c_{m}-P_{m})=(c_{n}-c_{m})-(P_{n}-P_{m}).$ (51)
Since $P_{n}-P_{m}$ is polynomial and bounded on the attractor, $v_{n}-v_{m}$ is bounded.
For invariants from different regularization classes, this algebraic synchronization reflects a structural property: although regularized variables follow different evolution equations, they remain coupled through the shared trajectory. The conservation laws enforce a coherence that prevents unbounded divergence.
IV.3 Physical Interpretation: Bounded History Accumulation
Each invariant $K_{n}$ defines a “measure of history” through $v_{n}(t)$, but this history does not accumulate indefinitely. It fluctuates within a bounded range determined by the attractor’s geometry.
From Eq. (49), the regularized variable at any instant is determined by the current state $(x,y,z)$ and the conserved value $c_{n}$. The evolution equation $v_{n}(t)=v_{n}(0)+\int_{0}^{t}Q_{n}\,ds$ and the algebraic representation $v_{n}=c_{n}-P_{n}$ are equivalent, but they emphasize different aspects. The former emphasizes history-dependence; the latter reveals that “history” is encoded in the instantaneous state.
This equivalence requires clarification to avoid logical ambiguity. While $v_{n}(t)=c_{n}-P_{n}(x,y,z)$ appears to render $v_{n}$ a function of the instantaneous state alone, this representation presupposes knowledge of the conserved value $c_{n}$. The quantity $c_{n}$ is determined by the initial condition: $c_{n}=P_{n}(x_{0},y_{0},z_{0})$ when $v_{n}(0)=0$. Crucially, $c_{n}$ cannot be recovered from the current state $(x(t),y(t),z(t))$ without knowledge of the initial state or, equivalently, the path integral $\int_{0}^{t}Q_{n}\,ds$. The “non-local” character of the constraint arises precisely because the instantaneous state does not determine the invariant value; the initial condition is required. In this sense, the constraint encodes information about the trajectory’s origin that is not contained in its current position.
Different invariants assign different “variations” to the same trajectory segment. When the trajectory passes through regions where $P_{n}$ changes rapidly, $v_{n}$ adjusts correspondingly. The rate of adjustment is encoded in $Q_{n}=\dot{v}_{n}$. The statistical analysis (Sec. VII) shows that these variations differ across classes: Class III invariants incur large-amplitude spikes in $Q_{\mathrm{III}}$ at lobe-switching events, while Class I invariants accumulate more smoothly.
The fluctuations of $v_{n}(t)$ trace a characteristic signature that encodes the trajectory’s relationship to the nullcline structure. The class structure partitions the invariants by dynamical sensitivity: Class III invariants respond to discrete topological transitions, while Class I invariants respond to continuous dynamics.
IV.4 Constraints on Trajectory Histories
Theorem 2 (Independence of History Constraints).
For any selection of three invariants $\{K_{\mathrm{I}},K_{\mathrm{II}},K_{\mathrm{III}}\}$ with one from each regularization class, the corresponding constraints are functionally independent:
@@ IV.4 Constraints on Trajectory Histories
provide three independent conditions that, together with the initial state $(x_{0},y_{0},z_{0})$, characterize a trajectory up to the discrete $\mathbb{Z}_{2}$ symmetry.
Proof.
Independence follows from the algebraic independence of the regularization polynomials. At $t=0$, with $v_{n}(0)=0$ (assuming $u_{n}(0)=0$), the invariant values are:
@@
Since $P_{\mathrm{I}}$, $P_{\mathrm{II}}$, and $P_{\mathrm{III}}$ are algebraically distinct polynomials, the map $(x_{0},y_{0},z_{0})\mapsto(c_{1},c_{2},c_{3})$ has full rank at generic points. Given $(c_{1},c_{2},c_{3})$, the polynomial equations $P_{n}(x,y,z)=c_{n}$ determine a discrete set of initial conditions.
For the canonical Triad $\{K_{5},K_{11},K_{13}\}$, the $3\times 3$ Jacobian matrix $\tilde{\mathbf{J}}$ of polynomial parts:
$\tilde{\mathbf{J}}=\begin{pmatrix}\partial_{x}P_{5}&\partial_{y}P_{5}&\partial_{z}P_{5}\\ \partial_{x}P_{11}&\partial_{y}P_{11}&\partial_{z}P_{11}\\ \partial_{x}P_{13}&\partial_{y}P_{13}&\partial_{z}P_{13}\end{pmatrix}$ (56)
has leading-order determinant for large $z$:
$\det(\tilde{\mathbf{J}})=\beta\sigma z^{2}(y-x)(y-x\rho)+\text{lower-order terms}.$ (57)
The full determinant is degree 5 in $(x,y,z)$; the zero locus is a lower-dimensional algebraic variety. ∎
Degeneracy locus. The determinant (57) vanishes on:
  1. 1.
    $z=0$: the horizontal plane through the origin;
  2. 2.
    $y=x$: the diagonal plane (Class I nullcline);
  3. 3.
    $y=x\rho$: a plane through the origin with slope $\rho$.
These loci have measure zero with respect to both Lebesgue measure and the SRB measure on the attractor. The fingerprint $(c_{1},c_{2},c_{3})$ provides a unique identifier at generic points.
IV.5 The Triad Construction
@@
provides a canonical set of three independent history constraints. The specific choice within each class affects only the numerical values, not the structural information.
Each Triad defines a “fingerprint” $(c_{1},c_{2},c_{3})$ that can be computed instantaneously from $(x_{0},y_{0},z_{0})$ via Eq. (55), since $v_{n}(0)=0$. The canonical Triad $\{K_{5},K_{11},K_{13}\}$ consists of even invariants under $\mathbb{Z}_{2}$, so the fingerprint identifies trajectories up to discrete symmetry.
@@ V.1 Invariants as Constraints on Trajectory Histories
The invariants $K_{n}=P_{n}(x,y,z)+v_{n}=c_{n}$ define constraints of a different character. The regularized variable $v_{n}(t)$ is a functional of the trajectory history:
@@
where $\gamma$ denotes the trajectory from $t=0$ to the current time. The constraint $K_{n}=c_{n}$ relates the instantaneous state $(x,y,z)$ to the accumulated history $v_{n}$; it is non-local in time.
Different invariants require different regularized variables because each imposes a distinct relationship between configuration and history. Invariants from different regularization classes “observe” the trajectory through different lenses: Class I through the yxy-x nullcline structure, Class III through the $z$-nullcline and its connection to lobe-switching events.
@@ V.2 Visualization of Constraint Surfaces
One can fix the regularized variable $v_{n}$ at a particular value and plot the resulting surface in $(x,y,z)$ space. Figure 1 displays such surfaces for the canonical Triad $\mathcal{T}=\{K_{5},K_{11},K_{13}\}$ along an unstable periodic orbit (UPO).
Several caveats apply:
  1. 1.
    The three surfaces correspond to different regularized variables $(v_{5},v_{11},v_{13})$, not a shared coordinate.
  2. 2.
    Along the UPO, regularized variables take specific values determined by integration over the orbit.
  3. 3.
    The surfaces intersect along the UPO because the invariant values were computed from this orbit’s initial conditions.
For any trajectory, the three constraints $K_{n}=c_{n}$ are satisfied simultaneously. Each constraint operates in its own extended space $(x,y,z,v_{n})$.
V.3 Resolution of the Structure-Chaos Relationship
The invariants illuminate the relationship between structure and chaos in the Lorenz system.
Structure: Each trajectory is characterized by its Triad fingerprint $(c_{1},c_{2},c_{3})$, determined instantaneously from initial conditions. The canonical Triad consists of even invariants, so the fingerprint identifies trajectories up to $\mathbb{Z}_{2}$ symmetry. Different trajectories (generically) have different fingerprints.
Chaos: Chaotic dynamics arise from sensitivity to perturbations in the fingerprint. Infinitesimally different initial conditions produce infinitesimally different fingerprints. While each fingerprint uniquely determines a trajectory, the mapping from fingerprint to long-time behavior inherits the Lyapunov instability of the underlying flow.
This resolves the apparent tension between “chaos” and “structure”: the invariants provide exact organization (a unique fingerprint per trajectory), while chaos persists as the exponential divergence of trajectories with nearby fingerprints.
V.4 Computational Complexity
The history-dependent constraints do not simplify the prediction of future states. Determining $(x(t),y(t),z(t))$ from $(x_{0},y_{0},z_{0})$ still requires numerical integration of the Lorenz equations. The regularized variable $v_{n}(t)=\int_{0}^{t}Q_{n}\,ds$ requires knowledge of the complete trajectory up to time $t$.
The computational complexity in the invariant formulation equals that of the original Lorenz system. The constraints provide structural insight and classification tools (Triad fingerprints), but do not reduce the dynamics to quadratures as classical integrability would.
V.5 Connection to Inertial Manifolds
The invariant structure resonates with, but differs from, the theory of inertial manifolds in dissipative systems [12]. An inertial manifold is a finite-dimensional, exponentially attracting, positively invariant smooth manifold that contains the global attractor.
The differences are significant:
  1. 1.
    Temporal validity: Inertial manifolds describe asymptotic behavior ($t\to\infty$). The invariant constraints hold exactly for all $t\geq 0$.
  2. 2.
    Mechanism: Inertial manifolds arise from spectral gaps in the linearized dynamics. Invariant constraints arise from the algebraic structure of the embedding.
  3. 3.
    Nature of constraint: Inertial manifolds are geometric objects in the instantaneous state space. Invariant constraints involve path functionals and are non-local in time.
V.6 Periodic Orbit Classification
Unstable periodic orbits (UPOs) form the “skeleton” of the strange attractor [13]. Their enumeration and classification are complicated by numerical instability and the lack of intrinsic invariant labels.
The Triad fingerprint provides a resolution. Each UPO corresponds to a specific triple $(c_{1},c_{2},c_{3})$. This serves as an intrinsic fingerprint that is independent of how the orbit is parameterized or which point is chosen as the initial condition. Two orbits with the same fingerprint are identical (up to $\mathbb{Z}_{2}$ symmetry); distinct orbits have different fingerprints.
This transforms UPO enumeration from a numerical search problem into a geometric cataloguing problem.
V.7 Poincaré Section Construction
The visualization of constraint surface intersections (Fig. 1b) requires a carefully constructed Poincaré section. A standard canonical section (e.g., $z=\mathrm{const}$) produces tangent or nearly degenerate curve intersections that obscure the independence of the three constraints. We therefore employ an optimized generalized section.
Reference point selection. The section plane $\Pi$ is centered at a point $\vec{P}_{\mathrm{opt}}$ on the unstable periodic orbit. Starting from the UPO initial condition $(x_{0},y_{0},z_{0})=(-1.31956,-0.38627,22.60542)$ with period $T\approx 1.55865$, the cutting point is defined at time
$t_{\mathrm{cut}}=T/4+0.0587\approx 0.4484,$ (60)
yielding $\vec{P}_{\mathrm{opt}}=\vec{X}(t_{\mathrm{cut}})$. This point serves as the origin of the local coordinate system on the section.
Normal vector optimization. The section plane is defined by a normal vector $\vec{n}(\theta,\phi)$ parameterized in spherical coordinates. The optimal orientation is determined by maximizing the minimum angular separation between projected constraint curves. For each candidate normal satisfying the transversality condition $|\vec{n}\cdot\vec{v}|>0.1$ (where $\vec{v}$ is the flow velocity), the tangent vectors of the projected level curves are computed as $\vec{t}_{i}=\vec{n}\times\nabla K_{i}$. The objective function
$\mathcal{J}(\theta,\phi)=\min_{i\neq j}\angle(\vec{t}_{i},\vec{t}_{j})$ (61)
is maximized using differential evolution, ensuring that the three constraint curves intersect with maximum angular aperture.
Local coordinates. The two-dimensional section employs orthonormal coordinates $(\eta,\xi)$ defined on $\Pi$. The basis vectors are constructed as
[…]
$\vec{V}$ $=\mathrm{Normalize}(\vec{n}\times\vec{U}),$ (63)
where $\hat{k}$ is the unit vector along $z$ (or $\hat{j}$ if singular). Any point $(\eta,\xi)$ in the section maps to three-dimensional space via
$\vec{X}_{3D}=\vec{P}_{\mathrm{opt}}+\eta\,\vec{U}+\xi\,\vec{V}.$ (64)
The coordinates $\eta$ and $\xi$ represent Euclidean displacements on the optimized plane, ensuring undistorted visualization of the constraint geometry.
Figure 1: Visualization of history-dependent constraints in $(x,y,z)$ space. (a) An unstable periodic orbit (dark curve) at standard parameters with constraint surfaces for Triad $\mathcal{T}=\{K_{5},K_{11},K_{13}\}$: Class I surface (blue), Class II surface (green), Class III surface (red). The surfaces represent $K_{n}=c_{n}$ at fixed $v_{n}$ values evaluated along the UPO. (b) Optimized Poincaré section through $\vec{P}_{\mathrm{opt}}$ on the UPO. The constraint surfaces intersect the plane as curves: $K_{5}=c_{1}$ (blue), $K_{11}=c_{2}$ (green), $K_{13}=c_{3}$ (red). The curves meet at a single point (the UPO crossing) with maximized angular separation, demonstrating the geometric independence of the three constraints. Coordinates $(\eta,\xi)$ are Euclidean displacements on the optimized section plane.

### v2

@@ III.3 Structural Origin of the Regularized Evolution Equation
A key technical point: the evolution equation for the regularized variable $v_{n}$ emerges as a function of $(x,y,z)$ alone, independent of $v_{n}$ itself. Consider an invariant of the form $K_{n}=P_{n}(x,y,z)+p(x,y,z)\cdot u_{n}$. The conservation requirement $dK_{n}/dt=0$ yields:
@@
Defining $v_{n}=p\cdot u_{n}$ and differentiating:
$\dot{v}_{n}=\dot{p}\cdot u_{n}+p\cdot\dot{u}_{n}=-\dot{P}_{n}.$ (14)
The $u_{n}$-dependent terms cancel exactly, yielding $\dot{v}_{n}=Q_{n}(x,y,z)\equiv-\dot{P}_{n}$, which depends only on the physical coordinates.
III.4 Class I Invariants: Permutations 1abc
@@
III.5 Class II Invariants: Permutations 2abc
@@
III.6 Class III Invariants: Permutations 3abc
@@
III.7 Null Class: Permutations 4abc
The six permutations beginning with 4 yield no valid invariants. Consider permutation 4123, corresponding to $(u,x,y,z)$. The orthogonality ansatz assigns
@@ III.7 Null Class: Permutations 4abc
Schwarz’s theorem for mixed partial derivatives requires $\partial_{yu}C=\partial_{uy}C$. From Eqs. (35) and (37):
@@
Equality requires $0=x$, a constraint that restricts $C$ to the hyperplane $x=0$, which is not invariant under the Lorenz flow. No smooth invariant satisfying ansatz (33) exists throughout phase space.
The impossibility of null-class invariants admits a physical interpretation in terms of causal hierarchy. In the valid classes (I–III), the auxiliary variable $u$ is driven by the physical coordinates through an evolution equation $\dot{u}=f(x,y,z)$ determined by the consistency conditions. The null class attempts the reverse: placing $u$ first in the permutation ordering would require the physical flow to be determined by the auxiliary dynamics. This violates the fundamental asymmetry between the “system” (physical coordinates) and the “memory” (auxiliary variable). In the language of Mori-Zwanzig theory, the memory kernel must depend on the resolved variables, not vice versa.
The complete classification is summarized in Table 1.
@@ IV.1 Boundedness of Regularized Variables
A consequence of the conservation laws is that regularized variables $v_{n}(t)$ remain bounded for all time. From conservation $K_{n}=P_{n}(x,y,z)+v_{n}=c_{n}$:
@@
Since $P_{n}$ is polynomial and the Lorenz attractor is compact, $v_{n}(t)$ is bounded.
@@
IV.2 Constraints on Trajectory Histories
For any selection of three invariants $\{K_{\mathrm{I}},K_{\mathrm{II}},K_{\mathrm{III}}\}$ with one from each regularization class, the corresponding constraints
@@ IV.2 Constraints on Trajectory Histories
are functionally independent and, together with the initial state $(x_{0},y_{0},z_{0})$, characterize a trajectory up to the discrete $\mathbb{Z}_{2}$ symmetry.
At $t=0$, with $v_{n}(0)=0$, the invariant values are determined by
@@
Since $P_{\mathrm{I}}$, $P_{\mathrm{II}}$, and $P_{\mathrm{III}}$ are algebraically distinct polynomials, the map $(x_{0},y_{0},z_{0})\mapsto(c_{1},c_{2},c_{3})$ has full rank at generic points.
IV.3 The Triad Construction
@@
provides a canonical set of three independent history constraints. Each Triad defines a “fingerprint” $(c_{1},c_{2},c_{3})$ that can be computed instantaneously from $(x_{0},y_{0},z_{0})$ via Eq. (45). The canonical Triad $\{K_{5},K_{11},K_{13}\}$ consists of even invariants under $\mathbb{Z}_{2}$, so the fingerprint identifies trajectories up to discrete symmetry.
@@ V.1 Invariants as Constraints on Trajectory Histories
The invariants $K_{n}=P_{n}(x,y,z)+v_{n}=c_{n}$ define constraints of a different character. The regularized variable $v_{n}(t)$ emerges as a functional of the trajectory history:
@@
where $\gamma$ denotes the trajectory from $t=0$ to the current time. The constraint $K_{n}=c_{n}$ couples the instantaneous state $(x,y,z)$ to the accumulated history $v_{n}$; it is non-local in time.
It is worth noting that invariants from different regularization classes “observe” the trajectory through fundamentally different lenses: Class I responds to the yxy-x nullcline structure, while Class III responds to the $z$-nullcline and its connection to lobe-switching events. This geometric distinction drives the statistical divergence observed in the chaotic regime.
@@ V.2 Visualization of Constraint Surfaces
Figure 1 displays constraint surfaces for the canonical Triad $\mathcal{T}=\{K_{5},K_{11},K_{13}\}$ anchored to an unstable periodic orbit (UPO). The three-dimensional visualization (panel a) depicts the level surfaces of the invariant polynomials in the Lorenz phase space: $P_{5}(x,y,z)=c_{1}$ (Class I, blue), $P_{11}(x,y,z)=c_{2}$ (Class II, green), and $P_{13}(x,y,z)=c_{3}$ (Class III, red), where the constants $c_{n}$ are determined by the UPO geometry. A representative chaotic trajectory illustrates the system’s evolution relative to these intersecting conserved manifolds.
Panel (b) presents a two-dimensional cross-section on an optimized Poincaré plane. The local coordinates $(\eta,\xi)$ are defined by $\mathbf{P}(\eta,\xi)=\mathbf{P}_{0}+\eta\,\mathbf{\hat{u}}+\xi\,\mathbf{\hat{w}}$, where $\mathbf{P}_{0}$ is a point on the unstable periodic orbit and $\{\mathbf{\hat{u}},\mathbf{\hat{w}}\}$ form an orthonormal basis of a plane whose orientation was numerically optimized to maximize the angular separation between the three level curves. The curves meet at a single point (the UPO crossing), confirming that all three constraints are simultaneously satisfied.
Figure 1: Geometric structure of the regularization classes. (a) Three-dimensional representation of the level surfaces of the invariant polynomials in the Lorenz phase space: $P_{5}=c_{1}$ (Class I, blue), $P_{11}=c_{2}$ (Class II, green), and $P_{13}=c_{3}$ (Class III, red), where the constants are anchored to an unstable periodic orbit. A representative chaotic trajectory (dark curve) illustrates the system’s evolution relative to these conserved manifolds. The black dot marks a point on the UPO. (b) Two-dimensional cross-section in local coordinates $(\eta,\xi)$ on an optimized Poincaré plane showing the topological relationship between the three regularization classes: $K_{5}=c_{1}$ (Class I, blue solid), $K_{11}=c_{2}$ (Class II, green dashed), and $K_{13}=c_{3}$ (Class III, red dot-dashed). The intersection point corresponds to configurations where all three polynomial constraints are satisfied simultaneously on the UPO. Parameters: $\sigma=10$, $\rho=28$, $\beta=8/3$.